A disaster relief team consists of engineers and doctors in the ratio of 2:5
  1. if there are 18 engineers, find the number of doctors.
  2. if there are 65 doctors, find the number of engineers.

Grade:7

2 Answers

Abhirup Roy
7 years ago
L
Umakant biswal
7 years ago
dear. 
so, as per the ratio . 
no of engg = 2x 
no of doctor = 5x 
then as given in ques , no of engg = 18 
so, 2x = 18 
so, x = 18/2 = 9 
so, no of doctor = 5x = 5*9 = 45 
in the 2nd case , no of doctor = 5x = 65 
so, x = 65/5 = 13
so, no of engg = 2x = 2*13 = 26 
HOPE IT CLEARS NOW 
REGARDS